# Graph Max Shift: A Hill-Climbing Method for Graph Clustering

### v1

@@ 3.2.3 Connection to Morse clustering
As we were preparing the manuscript, we realized that Graph Max Shift could be seen as a special case of a class of clustering algorithms called Morse Clustering recently proposed by Strazzeri and Sánchez-García [26]. Although the class of algorithms bears the name of ‘Morse’, the motivation for their introduction is entirely different from ours.
A Morse Clustering algorithm is based on an edge preorder and a node preorder, which we denote by $\preceq_{E}$ and $\preceq_{N}$, respectively. The choice of these preorders defines the algorithm. Indeed, starting from node $i$, if the path is at $i_{t-1}$ at step $t-1$, the path ends there unless there is a node $j\leftrightarrow i$ such that $(i,k)\preceq_{E}(i,j)$ for all $k\leftrightarrow i$; and $i\preceq_{N}j$ but $j\npreceq_{N}i$, in which case $i_{t}=j$ and the path continues. Nodes whose paths end at the same node are clustered together.
Graph Max Shift can be seen to belong in that class of algorithms. Assume that Graph Max Shift breaks ties between nodes by selecting the largest one when nodes are seen as an integers. Then define the node preorder $i\preceq_{N}j$ if $i=j$, or $q_{i}<q_{j}$, or $q_{i}=q_{j}$ and $i<j$; and define the edge preorder $(i,j)\preceq_{E}(i,k)$ if $j\preceq_{N}k$. Then it is a matter of checking that the Morse Clustering algorithm with these two preorders coincides with Graph Max Shift.
Although our algorithm belongs to the class of Morse Clustering algorithms, our motivation and that of Strazzeri and Sánchez-García [26] are very different. The impetus for proposing Morse Clustering algorithms is to address the impossibility result of Kleinberg [20]. In the context of weighted graphs, Kleinberg proposes three axioms that a clustering algorithm should satisfy, scale-invariance, richness, and consistency, and then proceeds to show that no algorithms satisfies all three. Strazzeri and Sánchez-García offer some criticism of the consistency axiom, propose to replace it by a weaker axiom which they call monotonic consistency, and show that there is at least one Morse Clustering algorithm which satisfies this new axiom, and the other two axioms of scale-invariance and richness.
Due to the fact that the intention is very different to ours in the present work, it is not surprising that there is no analysis in [26] of asymptotic consistency in the sense considered here, i.e., the recovery of gradient ascent flow clustering.
@@ 3.3 Consistency
In this section, we establish the asymptotic (statistical) consistency of Graph Max Shift when applied to $\mathcal{G}(\mathcal{Y},\epsilon_{n})$, when $\mathcal{Y}:=\{y_{1},\dots,y_{n}\}$ is generated iid from a density $f$ on $\mathbb{R}^{d}$ which satisfies some regularity conditions, and when $\epsilon_{n}\to 0$ slowly enough. In more detail, the reference partition is the one defined by Fukunaga and Hostetler [12], namely, $i$ and $j$ are in the same cluster if $y_{i}$ and $y_{j}$ belong to the basin of attraction of the same mode of $f$. We say that Graph Max Shift (or any other clustering algorithm) is consistent if the partition it outputs has Rand index [24] with respect to the reference partition tending to 1 in probability as $n\to\infty$, or more loosely, that the fraction of nodes that are incorrectly clustered converges to zero as the sample size increases without bound.
@@
The consistency of Graph Max Shift is not obtained as a special case of the consistency results established in [2], which crucially depend on the use of a second order kernel. The simplicity of Max Shift allows us to simplify the arguments in [2] (which were developed for generic hill-climbing methods) and, in particular, avoid the more complex aspects having to do with the stability of the basins of attraction to perturbation of the density.
@@
$\mathcal{Y}:=\{y_{1},\dots,y_{n}\}$ is generated iid from a density $f$ on $\mathbb{R}^{d}$ with compact support, twice continuously differentiable, and Morse. Then, with a large enough choice of tuning parameter $\tau$, Graph Max Shift applied to $\mathcal{G}(\mathcal{Y},\epsilon_{n})$ is consistent when $\epsilon_{n}\to 0$, and $\epsilon_{n}^{\max\{d+4,2d\}}n/\log n\to\infty$.
The proof of Theorem 3.3 is in Section 6. We note that the assumption of compact support is for simplicity, as the consistency result extends to densities that simply converge to zero at infinity as long as $f$ and its first and second derivatives are bounded.
We note that in practice, the choice of tuning parameter $\tau=1$ — meaning that clusters whose representative nodes are direct neighbors are merged — is often sufficient for good clustering performance, though without the merging post-processing step, it is possible that some of the population clusters defined by the density gradient ascent basins of attraction could be split into multiple clusters. However, without merging, Graph Max Shift applied to $\mathcal{G}(\mathcal{Y},\epsilon_{n})$ is weakly consistent when $\epsilon_{n}\to 0$ and $n\epsilon_{n}^{d+2}/\log n\to\infty$, in the sense that the fraction of pairs of points that belong to different population clusters but are clustered together by the algorithm tends to zero.
@@ 4 Numerical experiments
In this section, we demonstrate the behavior of Graph Max Shift on random geometric graphs with different underlying densities. The implementation is built upon the igraph package [10]. Python scripts to reproduce all figures in this section are available at https://github.com/lizzycoda/GraphMaxShift.
@@
In Figure 4.1 we demonstrate the consistency of Graph Max Shift on various (realizations of) random geometric graphs, where the underlying density $f$ is a Gaussian mixture with different numbers of components. We use the same mixture distributions as used in [7]. For each distribution, we sample $n=10^{4}$ data points, and apply the Graph Max Shift method with $\tau=1$. The value of $\epsilon$ used to construct $\mathcal{G}(\mathcal{Y};\epsilon)$ is indicated in each figure, and was chosen to obtain a low clustering error. The solid black lines represent the borders between the clusters defined by the gradient ascent flow and were found using the numerical method described in Section 3.1 of [7]. The method is based on Theorem 1 in [25]. For all distributions, most points are clustered together correctly, with some errors near the borders of the different basins of attractions or in low density regions.
@@
Figure 4.1: Graph Max Shift applied to $\mathcal{G}(\mathcal{Y};\epsilon)$ where the sample, of size $n=10^{4}$, is drawn from different normal mixtures. The solid black lines represent borders between the basins of attraction (population clusters) and data points are colored according to the clustering obtained via Graph Max Shift. Points that belong to clusters with fewer than $25$ points are colored in gray.
@@
Figure 4.2 depicts the effects of adjusting $\epsilon$ in $\mathcal{G}(\mathcal{Y};\epsilon)$ holding the tuning parameter $\tau$ fixed. We use a fixed a sample of $n=10^{4}$ points from a bimodal Gaussian mixture. In this mixture, the separation between the two components is small so that the clustering task is more difficult. We plot the clustering produced by Graph Max Shift with $\tau=1$ over a range of values of $\epsilon$. Considering the perspective in Section 3.2.1, observe that when $\epsilon$ is too small, the density estimator is not very smooth, which results in too many modes. In the clustering step, the small search radius obtains oversegmented clusters. When $\epsilon$ is large, the density estimator too smooth in the sense that some modes of the underlying density are not recovered, and coupled with the large search radius that allows points to ‘cross’ the basins of attraction, the result is an inaccurate clustering. Using the same sample, Figure 4.3, depicts the effects of adjusting the tuning parameter $\tau$ with $\epsilon$ fixed. When $\tau$ is too small the left cluster is split into multiple clusters, but by increasing $\tau=2$, the nodes in the left component are merged together and most points are correctly clustered. When $\tau$ is too large, clusters in different basins of attraction are merged together.
@@
Figure 4.2: Graph Max Shift applied to $\mathcal{G}(\mathcal{Y};\epsilon)$ with data drawn from a bimodal Gaussian mixture. The top row shows the obtained clustering with the indicated $\epsilon$ and $\tau=1$. The bottom row depicts the degree, which is proportional to the value of the density estimator implicitly computed. Additionally, each plot includes a ball of radius $\epsilon$ in the bottom right for reference.
Figure 4.3: Graph Max Shift applied to $\mathcal{G}(\mathcal{Y};\epsilon)$ with data drawn from a bimodal Gaussian mixture. The top row shows the obtained clustering with $\epsilon=0.06$ fixed as $\tau$ is varied in $\{1,2,3\}$. The nodes associated with the clusters are plotted in black. Additionally, each plot includes a ball of radius $\epsilon$ in the bottom right for reference.
@@
Finally, in Figure 4.4 we plot some paths constructed by Graph Max Shift in the setting of Figure 4.2 corresponding to $\epsilon=0.4$ and $\tau=1$. We also include an approximation of the Max Shift sequence $(x_{k})$ based on $f$ itself. The sequences closely follow one another and end close to the associated mode.
@@
Figure 4.4: Graph Max Shift path $(y_{k})$ in green and an approximation of the Max Shift path $(x_{k})$ based on $f$ itself in yellow. Modes of the density are plotted as stars.
@@
We find that for some densities, even when the data is just two dimensional, Graph Max Shift can have a large variance, even with sample sizes as large as $n\approx 10^{5}$.
@@ 5 Discussion
5.1 Merging parameter
As we could not discard the presence of multiple points of local maximum degree in the vicinity of the same density mode (as specified by Lemma 6.8), we opted for a post-processing step where clusters with nearby representative nodes are merged together. Even though less than elegant, this sort of step is not unusual in the literature. The choice of this tuning parameter is not straightforward, as is the case of most tuning parameters of clustering (and other unsupervised learning) methods, but is fairly intuitive here: set the parameter $\tau$ to less than half the minimum distance between two cluster centers that the user wants to be able to distinguish. Side information specific to an application could provide some insight into that. (But, again, we do not want to emphasize the practical nature of the proposed method.)
@@ 5.2 Multi-hop variant
$i_{0}=i;\qquad i_{t}\in\operatorname*{arg\,max}\{q_{j}:j\leftrightarrow_{m}i_{t-1}\},\quad t\geq 1,$ (5.1)
@@
The consistency result established in Theorem 3.3 generalizes under some conditions on $\epsilon$ and $m$ in a straightforward manner. Choosing $m$ is nontrivial, as is typically the case for the choice of tuning parameters in unsupervised learning. We also note that, in practice, unless $m$ is quite small, the algorithm can be slow to run.
@@ 5.3 Weighted graphs
We have focused on unweighted graphs, but the algorithm can be easily adapted to work with weighted graphs. Consider a graph with $n$ nodes, taken to be $[n]:=\{1,\dots,n\}$ without loss of generality, and dissimilarity $\delta_{ij}$ between nodes $i$ and $j$, with $\delta_{ij}=\infty$ if $i$ and $j$ are not neighbors. A straightforward way to apply Graph Max Shift to such a graph is to apply a threshold so that it becomes an unweighted graph. Specifically, $i$ and $j$ would be redefined as neighbors if $\delta_{ij}\leq h$, where $h$ is now a tuning parameter. The multi-hop variant described in Section 5.2 is best implemented by defining the local search area around node $i$ by $\{j:\delta_{ij}\leq r\}$, where $r\geq h$ is another tuning parameter. The correspondence with the number of hops $m$ introduced in Section 5.2 is roughly $r\approx mh$.
The situation most appropriate here would be a weighted geometric graph, $\mathcal{G}(\mathcal{Y},\epsilon)$, where the weights are given by $\delta_{ij}=\|y_{i}-y_{j}\|$ if $\|y_{i}-y_{j}\|\leq\epsilon$ and =absent=$\infty$= ∞ otherwise. The consistency result established in Theorem 3.3 generalizes in this setting under some conditions on $\epsilon$, $h$, and $r$, following the same line of arguments. Here too, choosing the tuning parameters $h$ and $r$ is challenging.

### v2

@@ 3.2.3 Connection to Morse clustering
Graph Max Shift can be seen as a special case of a class of clustering algorithms called Morse clustering recently proposed by Strazzeri and Sánchez-García (2022). A Morse clustering algorithm is based on an edge preorder and a node preorder, which we denote by $\preceq_{E}$ and $\preceq_{N}$, respectively. The choice of these preorders defines the algorithm, which proceeds as follows: Starting at $i_{0}=i$, if the path is at $i_{t-1}$ at step t1t-1, the path ends there unless 1) there is a node $j\leftrightarrow i_{t-1}$ such that $(i_{t-1},k)\prec_{E}(i_{t-1},j)$ for all $k\leftrightarrow i_{t-1},k\neq j$; and 2) $i_{t-1}\prec_{N}j$; if conditions 1) and 2) are met, then the node $j$ is unique, and the process continues with $i_{t}=j$. Doing this for all nodes $i$ defines what Strazzeri and Sánchez-García (2022) call the Morse flow. Nodes whose flows end at the same terminal node are clustered together.
Graph Max Shift is a Morse clustering algorithm. Indeed, for convenience, assume that Graph Max Shift breaks ties between nodes by selecting the largest one when nodes are represented by their indices. Then define the node preorder $i\preceq_{N}j$ if $i=j$, or $q_{i}=q_{j}$ and $i<j$, or $q_{i}<q_{j}$; and define the edge preorder $(i,j)\preceq_{E}(i,k)$ if $j\preceq_{N}k$. Then it is a matter of checking that the Morse clustering algorithm with these two preorders coincides with Graph Max Shift.
Although our algorithm belongs to the class of Morse clustering algorithms, our motivation and that of Strazzeri and Sánchez-García (2022) are very different. Indeed, the impetus in Strazzeri and Sánchez-García (2022) for proposing Morse clustering algorithms is to address the impossibility result of Kleinberg (2002). In the context of weighted graphs, Kleinberg proposes three axioms that a clustering algorithm should satisfy, scale-invariance, richness, and consistency, and then proceeds to show that no algorithm satisfies all three. Strazzeri and Sánchez-García (2022) offer some criticism of the consistency axiom, propose to replace it by a weaker axiom which they call monotonic consistency, and show that there is at least one Morse clustering algorithm which satisfies this new axiom, and the other two axioms of scale-invariance and richness.
Due to the fact that the intention is very different to ours in the present work, it is not surprising that there is no analysis in Strazzeri and Sánchez-García (2022) of asymptotic consistency in the sense considered here, i.e., the recovery of gradient ascent flow clustering in the context of a random geometric graph.
@@ 3.3 Consistency
In this section, we establish the asymptotic (statistical) consistency of Graph Max Shift when applied to $\mathcal{G}(\mathcal{Y},\epsilon_{n})$, when $\mathcal{Y}:=\{y_{1},\dots,y_{n}\}$ is generated iid from a density $f$ on $\mathbb{R}^{d}$ which satisfies some regularity conditions, and when $\epsilon_{n}\to 0$ slowly enough. In more detail, the reference or population partition is the one defined by Fukunaga and Hostetler (1975), namely, $i$ and $j$ are in the same population cluster if $y_{i}$ and $y_{j}$ belong to the basin of attraction of the same mode of $f$. Consider the distance between two partitions of $[n]$ defined as the fraction of pairs of nodes $i\neq j$ that are grouped together in one partition but not in the other partition. This distance is equal to one minus the Rand index Rand (1971) — a measure of similarity between partitions popular in the clustering literature. We say that Graph Max Shift, or any other clustering algorithm for that matter, is consistent if the distance between the partition it outputs and the reference partition tends to zero in probability as $n\to\infty$, or equivalently, if the Rand index tends to one. By contrast, a clustering method is weakly consistent if the fraction of pairs of points that belong to different population clusters and are clustered together by the method tends to zero in probability. In essence, a method is consistent if it recovers the population partition asymptotically, while a method is weakly consistent if it recovers a possibly finer partition asymptotically.
More formally, and for future reference in the proof section, let $(\mathbf{B}_{k})$ denote the basins of attraction of density modes. (Note that there could well be infinitely many of them, although only countably many under our assumptions on the density.) Given $\mathcal{Y}$, the reference clusters are $\mathbf{I}_{k}=\{i:y_{i}\in\mathbf{B}_{k}\}$. Under our assumptions on the density, with probability one, $n]=\bigcup_{k}\mathbf{I}_{k}$, and this is the reference partition of the graph in the present context. A cluster method takes the graph adjacency matrix $A$ and returns a partition of $[n]$, say $n]=\bigcup_{k}\widehat{\mathbf{I}}_{k}$. We write $\mathbf{I}(i,j)=1$ if there is $k$ such that $i,j\in\mathbf{I}_{k}$, and $\mathbf{I}(i,j)=0$ otherwise, and $\widehat{\mathbf{I}}(i,j)$ has a similar meaning. Under the model of a random geometric graph described above, the method is weakly consistent if
$\frac{\#\big\{i<j:\widehat{\mathbf{I}}(i,j)<\mathbf{I}(i,j)\big\}}{\binom{n}{2}}\longrightarrow 0,\ \text{in probability as $n\to\infty$};$ (3.6)
and it is consistent if
$\frac{\#\big\{i<j:\widehat{\mathbf{I}}(i,j)\neq\mathbf{I}(i,j)\big\}}{\binom{n}{2}}\longrightarrow 0,\ \text{in probability as $n\to\infty$}.$ (3.7)
(The proportion in (3.7) is the distance between the partitions $(\mathbf{I}_{k})$ and $(\widehat{\mathbf{I}}_{k})$.)
@@
The consistency of Graph Max Shift is not obtained as a special case of the consistency results established in Arias-Castro and Qiao (2025), which crucially depend on the use of a second order kernel. The simplicity of Max Shift allows us to simplify the arguments in Arias-Castro and Qiao (2025) (which were developed for generic hill-climbing methods) and, in particular, avoid the more complex aspects having to do with the stability of the basins of attraction to perturbations of the density.
@@
Suppose $\mathcal{Y}_{n}:=\{y_{1},\dots,y_{n}\}$ is generated iid from a density $f$ on $\mathbb{R}^{d}$ with compact support, twice continuously differentiable, and of Morse regularity, and consider Graph Max Shift applied to the adjacency matrix of $\mathcal{G}(\mathcal{Y}_{n},\epsilon_{n})$ with $\epsilon_{n}\to 0$.
  1. (i)
    Without merging. If $\epsilon_{n}\gg\big(\frac{\log n}{n}\big)^{1/(d+2)}$, the method is weakly consistent.
  2. (ii)
    With merging. There is $C=C(f)$ such that, if $\tau\geq 1/C$ and $\epsilon_{n}\gg\big(\frac{\log n}{n}\big)^{1/\max\{d+4,\,2d\}}$, the method is consistent.
The proof of Theorem 3.3 is in Section 6. We note that the assumption of compact support is for simplicity, as we anticipate the consistency result extends to densities that simply converge to zero at infinity as long as $f$ and its first and second derivatives are bounded — which is the setting considered in Arias-Castro and Qiao (2025). The conditions on $\epsilon_{n}$ are not particularly intuitive and arise in the course of the proof. In particular, ignoring the logarithmic factor, we do not know if the exponent is optimal in either (i) or (ii) — although this sort of ‘curse-of-dimensionality’ dependence on the dimension is to be expected in the nonparametric setting that we consider given where only smoothness assumptions are made on the density.
The lower bound on the number of hops $\tau$ for merging in post-processing depends on $f$ in particular through the distance between its modes. We note that in practice, the choice of tuning parameter $\tau=1$ — meaning that clusters whose representative nodes are direct neighbors are merged — is often sufficient for good clustering performance, though without the merging post-processing step, it is possible that some of the population clusters defined by the density gradient ascent basins of attraction could be split into multiple clusters.
@@ 4 Numerical Experiments
In this section, we demonstrate the behavior of Graph Max Shift on random geometric graphs with different underlying densities. The implementation is built upon the igraph package Csardi and Nepusz (2006). Python scripts to reproduce all figures in this section are available at https://github.com/lizzycoda/GraphMaxShift.
@@
In Figure 4.1 we demonstrate the consistency of Graph Max Shift on various (realizations of) random geometric graphs, where the underlying density $f$ is a Gaussian mixture with different numbers of components. We use the same mixture distributions as used in Chacón (2015). For each distribution, we sample $n=10^{4}$ data points, and apply the Graph Max Shift method without merging ($\tau=1$). The value of $\epsilon$ used to construct $\mathcal{G}(\mathcal{Y};\epsilon)$ is indicated in each figure, and was chosen to obtain a low clustering error. The solid black lines represent the borders between the clusters defined by the gradient ascent flow and were found using the numerical method described in Section 3.1 of Chacón (2015). The method is based on Theorem 1 in Ray and Lindsay (2005). For all distributions, most points are clustered together correctly, with some errors near the borders of the different basins of attractions or in low density regions.
@@
Figure 4.1: Graph Max Shift applied to $\mathcal{G}(\mathcal{Y};\epsilon)$ where the sample, of size $n=10^{4}$, is drawn from different normal mixtures. The solid black lines represent borders between the basins of attraction (population clusters) and the triangles represent the locations of the modes of the density. As the means of the mixture components all either coincide with the modes or are very close to the modes, we do not plot them. Data points are colored according to the clustering obtained via Graph Max Shift. Points that belong to clusters with fewer than $25$ points are colored in gray.
@@
Figure 4.2 depicts the qualitative effects of adjusting $\epsilon$ in $\mathcal{G}(\mathcal{Y};\epsilon)$ holding the tuning parameter $\tau$ fixed. When $\epsilon$ is too small, the density estimator is not very smooth, which results in too many modes. In the clustering step, the small search radius obtains oversegmented clusters. When $\epsilon$ is large, the density estimator is too smooth in the sense that some modes of the underlying density are not recovered, and coupled with the large search radius that allows points to ‘cross’ the basins of attraction, the result is an inaccurate clustering.
In Figure 4.3 we quantify the effects of adjusting $\epsilon$, and plot the weak clustering error as defined in (3.6) and the clustering error as defined in (3.7). In the lefthand plot, the weak clustering error remains small over a range of $\epsilon$ small enough. Note that when $\epsilon$ is smaller than the minimum distance between two points, each vertex will correspond to its own cluster and the weak clustering error will be zero. In contrast, the condition on $\epsilon$ in Theorem 3.3 is imposed to ensure that points in the basin of attraction of a particular mode are moved by Graph Max Shift to a point close to that mode. For example, in Figure 4.2 when $\epsilon=0.18$ or $\epsilon=0.2$ the weak clustering error is small, though not all terminal nodes are close to the modes of the density. When $\epsilon=0.3$, the weak clustering error remains small, and the terminal nodes are all close to the modes of the density. In the righthand plot of Figure 4.3, the clustering error is small for a range of $\epsilon$ not too small and not too large. This reflects the oversegmentation when $\epsilon$ is too small and the crossing of the basins of attraction when $\epsilon$ is too large. Moreover, as $n$ increases, the values of $\epsilon$ for which the error is small moves to the left, reflecting the condition in Theorem 3.3.
@@
Figure 4.2: Graph Max Shift applied to $\mathcal{G}(\mathcal{Y};\epsilon)$ with data drawn from a trimodal Gaussian mixture. The top row shows the obtained clustering with the indicated $\epsilon$ and $\tau=1$. Terminal nodes are highlighted in black. The bottom row depicts the degree of each node, which is proportional to the value of the density estimator implicitly computed. Additionally, each plot includes a ball of radius $\epsilon$ in the bottom right for reference. The errors of each clustering are quantified in Figure 4.3
Figure 4.3: The error as a function of $\epsilon$ when Graph Max Shift is applied to $\mathcal{G}(\mathcal{Y};\epsilon)$ with $\tau=1$ for data drawn from the trimodal Gaussian mixture for different values of $n$. The left plot shows the weak clustering error (3.6) and the right plot shows the clustering error (3.7). Solid lines indicate the error averaged over $100$ simulations, and the shaded regions indicate empirical 80% intervals. Clusterings for select values of $\epsilon$ when $n=10^{4}$ are shown in Figure 4.2.
@@
Finally, in Figure 4.4 we plot some paths constructed by Graph Max Shift for the trimodal mixture corresponding to $\epsilon=0.4$ and $\tau=1$. We also include an approximation of the Max Shift sequence $(x_{k})$ based on $f$ itself. The proof of the main theorem relies on the sequences closely following one another, and indeed we observe this and see that they end close to the associated mode.
@@
Figure 4.4: Graph Max Shift path $(y^{k})$ in green and an approximation of the Max Shift path $(x_{k})$ based on $f$ itself in navy blue. Modes of the density are plotted as triangles.
@@
We find that for some densities, even when the data is just two dimensional, Graph Max Shift can have a large variance, even with sample sizes as large as $n\approx 10^{5}$. Additional quantitative plots for the other Gaussian mixtures in Figure 4.1 are included in the Appendix which also depict this variance. We also include an additional example where the clustering task is more difficult and adjusting the tuning parameter is necessary to achieve small clustering error.
@@ 5.1 Multi-hop variant
$i_{0}=i;\hskip 18.49988pti_{t}\in\operatorname*{arg\,max}\{q_{j}:j\leftrightarrow_{m}i_{t-1}\},\qquad t\geq 1,$ (5.1)
@@
We anticipate the consistency result established in Theorem 3.3 to generalize in a straightforward manner under some appropriate conditions on $\epsilon$ and $m$. We note that choosing $m$ is nontrivial, as is typically the case for the choice of tuning parameters in unsupervised learning. We also note that, in practice, unless $m$ is quite small, the algorithm can be slow to run.
@@ 5.2 Weighted graphs
We have focused on unweighted graphs, but the algorithm can be easily adapted to work with weighted graphs. Consider a graph with $n$ nodes, taken to be $n]:=\{1,\dots,n\}$ without loss of generality, and dissimilarity $\delta_{ij}$ between nodes $i$ and $j$, with $\delta_{ij}=\infty$ if $i$ and $j$ are not neighbors. A straightforward way to apply Graph Max Shift to such a graph is to apply a threshold so that it becomes an unweighted graph. Specifically, $i$ and $j$ would be redefined as neighbors if $\delta_{ij}\leq h$, where $h$ is now a tuning parameter. The multi-hop variant described in Section 5.1 is best implemented by defining the local search area around node $i$ as $\{j:\delta_{ij}\leq r\}$, where $r\geq h$ is another tuning parameter. The correspondence with the number of hops $m$ introduced in Section 5.1 is roughly $r\approx mh$ — see Arias-Castro et al. (2021), where this simple observation plays a central role. The most appropriate situation here would be a weighted geometric graph, $\mathcal{G}(\mathcal{Y},\epsilon)$, where the weights are given by $\delta_{ij}=\|y_{i}-y_{j}\|$ if $\|y_{i}-y_{j}\|\leq\epsilon$ and $\infty$ otherwise.
We expect the consistency result established in Theorem 3.3 to generalize to this setting under some appropriate conditions on $\epsilon$, $h$, and $r$, following the same line of arguments. Here too, choosing the tuning parameters $h$ and $r$ is challenging.